aluminum has a specific heat of 0.910 j/gc if 15.2 g of al are heated to a temperature of 99.5 c and then immersed in 50.0 g of water what is the final temperature of the water the water is initially at 25 c
Answer #1

heat lost by Al + heat gain by coate (ms AT) HO CmsaTA (IS.29xa91T-99.5)) J - 25)) =o +(sogx u1 on solving, T = 24.6